A Complete Bibliography of Publications by, and about, C. W. Clenshaw

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

19 August 2024 Version 1.02

Abstract

This bibliography records publications of C. W. Clenshaw.

Title word cross-reference

6 × 9 [Cle67]. $7.50 [Cle63]. _{1}F_{1}(1; c; x) [dB76]. A^{∗}Φ(Bx) [Dun73a]. cos(iπ/n) [Oli75]. du [GS48]. Erfc(x) [Ric64]. Γ(x) [Ric64]. J_{0}(x) [WBR82]. J_{1}(x) [WBR82]. J_{n}(z) [CM83]. L^{2} [Pat76]. L_{∞} [Ric64]. Lg [JS69]. log n [KR79]. X [CN78, Sma85]. Y_{0}(x) [WBR82]. Y_{1}(x) [WBR82].

-spline [CN78]. -splines [JS69, Sma85].

1 [W.63]. 1966. [Cle67]. 1987 [Tur89]. 1992 [Ano92].

424 [Ged79, Gen72a].

absence [Car72b]. Academic [Cle72]. accelerate [Boy23]. Accuracy [SV73, Sut19a]. ACM [Gen72a]. adaptive [Maj21b, Oli72]. admitting [Bra72]. afford [CMM16]. Algebraic [Wan92, DGK13, Wan18]. Algorithm [Bar13, Gen72a, ATAAK17, AGR93, Bar98, Car72a, Car72b, CO80, Ell68, Has01, HNU14, Mat72, Smo02, dGR95, Ged79, Goo73]. algorithme [Car72a, Car72b]. Algorithms [CMW63, Bar02, CO86, Deu76, KEKT21]. Allan [CLG^{+}03]. almost [Ell73]. Alternative [Ada87]. analogue [AGR93].

Analysis [Ano78, CO51, Cle63, Cle78, Oli77, Sut19c, Tur89, XCWB13, Bar98, Ell68, Has01, JHC^{+}14, Mas70, New74, Olv04, Pry81, Sut19a, Cle68a]. analytic [FLR93]. Angenäherte [Fil64]. Application [Orl77a, CN78, HTN83, KKY07, Maj19, Oli80]. applications [Pie87, XCWB11, XCWB13]. approximant [Bak74]. Approximants [MC05b, MC05a]. Approximate [PDP12a, Cle63]. approximates [SH21]. approximating [PDP12b, Rou70]. Approximation [BdBSW66, Fil64, HF11, MP72, RT72, Tay69, Bas73, BCG74, Coh73, DK71, Dun73a, Dun73b, Dun75, FGR76, Ful77, HW77, Hol69, JHC^{+}14, KEKT21, Mas70, Mur71, NTT15, RR71, Row69, Sti65, Tal76, Tay68, TW74, Cle67, Cle66a]. Approximations [Cle54, Cle64, RP63, Sch79, And73, CL74, Cle74, Ell73, Ful77, Mat72, Neu74, Pat76, ST69, Tay70, WBR82, Cle72]. April [CLG^{+}03]. Arithmetic [CO87, COT89, CT89, LT92]. arrays [Ric64]. art [PTVF92a]. Arthur [Cle63]. Assessing [LRD19]. assurant [Car72a]. Automatic [CO51, CC60, HTN83]. axe [FGR76]. axis [FGR76].

B [Cle68a, Tod67]. Backward [Smo02]. band [JHC^{+}14]. based [MV97, SHN20, Smi65]. Berkeley [Wan92]. Bernstein [HW77]. Bessel [XCWB13, CP66, CM83, GCP67, LPM81, PB83, WBR82, XCWB11]. Best [Cle64, Mur71, Ell73, MP72, Mat72]. Better [Tre08]. Beyond [CO84]. bivariate [GC24]. blended [BCG74]. Blending [PDP12a, PDP12b]. Book [Cle63, Cle65, Cle66a, Cle67, Cle68a, CMM16]. Books [Cle72, W.63]. bound [Bar03]. boundary [CL23, EJJ08, Ura69]. Bounds [FLR93, Kam71a, Olv67, Bar02, Rou70, Smi82, XHC15].

C [W.63]. C. [LE72]. Calculation [Cle63, CL23]. California [Wan92]. can [CMM16]. Cas [MP72]. case [CE60, MP72]. Cauer [Tod67]. Certain [Bak74, Deu76, Ell68, EJJ08, Hun70, Kam71b, TE72]. Charles [Olv04]. CHEBYLAN [LM76]. Chebyshev [Bas73, BS71, Cle55, Cle57, Cle60b, Cle62a, CN63, Cle64, CP66, Cle66b, CL74, CM83, DK71, Dun73a, Dun73b, Dun75, EL00, EL06, Fil64, GCP67, Han73, HTN83, Has19, Hay73, Jay74, JHC^{+}14, LM76, Luk76, MV97, MC05b, MC05a, Mur71, Oli69, Oli77, RP63, RR71, Row69, Sch78, Sut19b, TW74, Tod67, Ura69, W.63, WBR82, Cle67, Cle68a]. chemistry [SAM08]. circular [Sch78]. class [Dom14]. classes [TE72]. Clenshaw [Bra73, CL14, DGS11, DGK13, Dom14, Fil64, GI17, GC24, Goo73, KKR03, Loc69, Maj17, Maj19, Maj21b, Maj21a, MFA22, W.63, XCWB13, XHC15, AZK95, Ada87, AN91, AF84, ATAAK17, AGR93, Bar13, Bar98, Bar02, Boy23, BP75, Bra92, CE09, Cha68, CL23, DD11, Den86, ES88, EJJ08, EL00, EL06, EHM81, FLR93, Fil64, FW66, Ged79, Gen72a, Gen72b, Gen72c, GW22, GW23, Håv69, HST93, Has01, HS15, HF11, HNU14, HFWW22, Hun70, HS05, Imh63, JKH^{+}17, Jay74, JHC^{+}14, Kam71a, KEKT21, KKY07, Kus72, Küt84, LE72, LX19, Loc69, LRD19, Ma15, Mar86, MV97, MC05b, MC05a, NF03, NTT15, Not16, OS68, Oli71, Oli72, Oli77, Olv04, PDP12b, PDP12a, Pet94, PB83, Pie87, PTVF92b, RV07, Rei77, Rie71, RJ71, RJ72, SAM08, SHN20, SS78a, SS78b]. Clenshaw [SS80, Smi65, Smi82, SH11, Smi15, Smo02, SVZ08, Som13, SH21, SH22, Sut19a, Sut19b, Tre08, Tua73, VR11, Wal06, Wan18, WT07, XCWB11, XB12, XB13, XHW14, XrL11, vWGB19]. Cliffs [Cle67]. cm. [Cle72, W.63]. Coefficients [TE72, Rou70]. collocation [HFWW22, KKY07, LRD19]. Company [Cle63]. Comparison [Cle64, NTT15]. complex [CM83, DK71, Ma15, SV73]. Computation [AZK95, NF03, RV07, VR11, Wan92, Pat76, PB83, Pie87, Ura69]. computational [LRD19]. computations [Boy23]. Computer [Cle60a, CC60]. Computing [CO51, CGM^{+}61, PTVF92a, Gen72c]. Condition [Gau79, Car72b]. conditionnement [MP72]. conformal [HT08]. conjecture [LE72]. constants [Pet94]. Construction [Wal06, Som13]. containing [LPM81]. Contribution [Ano78]. control [SHN20]. Convergence [And73, Smi15, XB13, dB76, Bak74, Car72a, CS78, SS78a, Wan18, XB12, SS78b]. converting [Sal73]. corrected [HS15]. cosine [AZK95, Gen72c]. Cote [DD11]. cubature [SVZ08]. cubes [NR96]. Cubic [BdBSW66, CN78]. Curtis [XCWB11, XCWB13, Ada87, AN91, AF84, ATAAK17, Bar13, BP75, Bra73, Bra92, CE09, Cha68, CL14, CL23, DD11, Den86, DGS11, DGK13, Dom14, EJJ08, EHM81, FLR93, Fil64, FW66, GI17, GC24, Ged79, Gen72a, Gen72b, Gen72c, Goo73, Håv69, HST93, HS15, HF11, HNU14, HFWW22, HS05, Imh63, Jay74, Kam71a, KKR03, KEKT21, KKY07, Kus72, Küt84, LX19, Loc69, LRD19, Ma15, Maj17, Maj19, Maj21b, Maj21a, MFA22, Mar86, MV97, NTT15, Not16, OS68, Oli71, Oli72, PDP12b, PDP12a, Pet94, PB83, Pie87, Rie71, RJ71, RJ72, SHN20, SS78a, SS78b, SS80, Smi65, Smi82, SH11, Smi15, SVZ08, Som13, SH21, SH22, Sut19a, Sut19b, Tre08, Wal06, Wan18, WT07, XB12, XB13, XHW14, XHC15, XrL11, vWGB19]. Curtis-rational [XrL11]. Curtis-type [LX19, SH22]. Curve [Cle60a, CH65, Cle68b].

D [W.63]. D1 [Gen72a, Goo73]. Data [Ran77, LRD19]. December [CLG^{+}03]. defined [Hol69]. Definite [PDP12a, PDP12b]. definiteness [AF84].

Degree [HW77, Sma85, Tal76]. Densité [Car72a]. Density [Car72a]. derivative [Coh73]. derivatives [Has01, HF11, RT72]. Descriptions [Cle72, W.63]. deviation [Row69]. d’exponentielles [MP72]. difference [Olv67]. Differential [CO51, Tua73, BS71, Cle57, Cle60b, CN63, Oli69, Orl77a, Orl77b, SAM08]. Digital [Cle60a]. dimensional [Hay73, NR96]. direct [Ell73, SHN20]. discontinuous [KKR03]. discrete [AZK95]. Discussions [CO51]. distributed [dGR95]. double [Bas73]. doubly [Oli72]. doubly-adaptive [Oli72]. Dr. [Ano78]. d’une [MP72].

economized [CO55]. edited [Cle66a]. effect [Pat76]. efficient [KEKT21]. electromagnetics [LRD19]. element [EJJ08]. Elementary [Cle54]. elimination [Orl77b]. Enclosing [Has19]. endpoint [Wan18]. Englewood [Cle67]. ensuring [Car72a]. equation [CE60, Cle66b]. Equations [CO51, Cle65, Tua73, BS71, CL14, CL23, Cle57, Cle60b, CN63, EL00, EL06, EHM81, HFWW22, KKR03, Oli69, Olv67, Orl77a, Orl77b, SAM08, XrL11]. Erratum [XCWB13]. Error [Bra92, Cha68, Cle78, Ell68, Has01, HS15, Jay74, Kam71a, Kam71b, New74, OS68, Oli77, Rie71, RJ71, RJ72, Bar02, Bar03, Bra73, DGS11, FLR93, Küt84, Not16, Oli69, Pet94, Pry81, SS80, Smi82, SH11, Ura69, XHC15, Loc69]. errors [Pat76]. Erweiterung [Peh78]. estimate [HS15]. Estimates [Cha68, DGS11, Jay74, Küt84, Loc69, Rie71, RJ71, RJ72, SS80]. estimation [Bra73, Bra92, OS68, Oli69, Ura69]. Étude [Car72b]. Euler [Smi65]. Evaluating [Oli77, Has01, Hun70, Oli75]. Evaluation [KR79, PDP12a, Bar02, Bar03, CE09, LPM81, New74, Oli80, SH11]. Existence [TW74]. expansion [Luk76, Sal73]. Expansions [Cle64, Has19, MC05b, CM83, Han73, Hol69, MC05a, TE72]. experience [Gen72b]. experimental [Ful77]. Experiments [Sut19c]. explicit [Orl77a]. Exponential [CLOT86, Sch79, CO80, Luk76, Maj19, Ran77]. exponentials [MP72]. extended [Ada87, CS78, HST93, Maj21b]. Extension [Tua73, BP75, Hay73, Peh78]. Extensions [SH22].

Faber [ES88]. families [Dun73b, RR71]. Fast [JHC^{+}14, RV07, Som13, Wal06, XHW14]. Fehlerabschätzung [Bra73]. Fehlerabschätzungen [Loc69]. Fejer [DD11, FLR93, Som13, Wal06, WT07, XHW14]. Fejér-type [XHW14]. fewer [Pry81]. Filon [XCWB13, DGS11, DGK13, Dom14, GI17, GC24, Maj17, Maj19, Maj21b, Maj21a, MFA22, XCWB11, XHC15]. Filon-type [XCWB13, XCWB11]. finite [Ell68, Hun70]. first [Coh73]. fits [Ran77]. Fitting [Cle60a, CJ74, CH65, Cle68b]. flair [Olv04]. Floating [CO84]. fonction [MP72]. Form [Gau79]. formal [Hol69]. Formula [Håv69, PTVF92b, RV07, VR11, AZK95, Den86, HS05, JKH^{+}17, Mar86, NF03, OS68, SHN20, Smi65, Smi82, SH11, Smi15, SH21]. formulae [AF84, Not16]. Formulas [Kam71a, HT08, Jay74, Kam71b, Maj17]. formulation [Ada87]. Forsythe [Bar02]. FORTRAN [PTVF92a]. Forward [VR11, NF03]. Four [MC05b, MV97, MC05a]. Fourier [KEKT21, Oli77]. Fox [Ano92, Cle68a, Cle92a, Cle92b]. Fractional [GCP67, CP66, HF11, SHN20]. French [Car72a, Car72b, FGR76, MP72]. Function [Fil64, Boy23, CO80, Coh73, CM83, DK71, HS05, Kus72, LPM81, MP72, PB83, SH11, Smi15]. Functions [Cle54, Cle62a, CMW63, Cle72, CLOT86, GCP67, XB13, Bra72, CP66, Cle74, Deu76, FLR93, HW77, Hol69, NR96, Orl77a, Ran77, RR71, Som13, Tay69, TE72, WBR82, XB12, W.63, Cle66a].

Garabedian [Cle66a]. Gauss

[EJJ08, Jay74, Kam71b, PDP12a, Tre08, XB12, XB13]. Gaussian [Boy23]. Gegenbauer [HS05, SH11, Smi15]. general [Som13]. Generalisation [ES88]. generalization [GI17]. Generalized [CLOT86, HTN83, KKY07, Lin76]. German [Bra73, Fil64, Loc69, Peh78]. given [Tur89]. Global [Smi82]. Graph [GW22, GW23]. grid [Sut19b]. grids [NTT15]. Grunwald [CS78]. Grunwald-type [CS78].

H [Ano78, Cle63, Cle66a]. Haar [Car72b]. Hall [Cle67]. Hammerstein [EL00, EL06]. Hardy [Cle04]. having [ST69, Tay68, Tay69]. Her [W.63]. Hermite [Boy23]. High [NR96, Pet94]. High-dimensional [NR96]. highly [DGS11, DGK13, Dom14, GI17, GC24, KEKT21, LX19, Maj17, XCWB11, XCWB13, XHC15]. highly-oscillatory [Dom14]. Hopf [XrL11]. Hybrid [GF92]. hypersingular [LX19]. hypotheses [Car72a, Car72a].

Identification [DD11]. II

[Cle72, Gen72c, MC05b, MC05a, ST69]. III [Tay69]. IMA [XCWB13]. Image [JKH^{+}17]. implementation [AN91, SS80, XHW14]. Implementing [Gen72b, Gen72c, Ma15]. implicit [Orl77a]. Index [CO87, COT89, CT88, CT89, LT92]. Inequalities [OR78, Tod67]. infinite [SH21, SH22]. input [LRD19]. integral [And73, CL14, CL23, EHM81, HFWW22, KKR03, Ma15, Pie87]. Integrals [PDP12a, Ada87, CL23, DGS11, DGK13, Dom14, EJJ08, GI17, GC24, KEKT21, LPM81, LX19, Luk76, Maj17, Maj19, Maj21a, PDP12b, PB83, SH21, SH22, Wan18, XHC15, Cle63]. Integration [MV97, Sut19b, AN91, BS71, CE09, CC60, GF92, HF11, Imh63, LPM81, NR96, Pie87, SS78a, SS78b, SS80, Smi65, Cle65]. International [Wan92]. interpolants [BCG74]. interpolating [TW74]. interpolation [CN78, De 75, HTN83, SV73]. interpolatory [Küt84]. interval [HW77, SH21, SH22]. Introductory [COT89]. invariance [Bak74]. Inverse [RV07, NF03]. Iterative [EHM81]. iv [W.63].

J [Ano78]. James [Cle04]. jet [CE60]. Jordan [Tod67]. Journal [XCWB13]. July [Wan92].

Karup [De 75]. kernels [CL14, KKR03, LX19]. kind [CS78, Jai74, XrL11]. Kinds [MC05b, MV97, MC05a]. kink [WT07]. Krawtchouk [JKH^{+}17, RV07, VR11]. Krylov [Cle63].

L [Cle68a, Cle72, W.63]. Laboratory [W.63]. laminar [CE60]. Lancaster [Tur89]. language [LM76]. Laplace [HF11]. Laurent [MC05b, MC05a]. law [OR78]. least [KKY07]. least-squares [KKY07]. Lectures [Tur89]. Legendre [EJJ08, PDP12a]. Leja [LRD19, NTT15]. Leslie [Cle92b, Ano92, Cle92a]. Level [CO87, COT89, CT88, CT89, LT92]. Level-Index [CO87, COT89, CT88, CT89]. Light [CLG^{+}03]. Limited [XB13, XB12]. Linear [Has19, Tua73, Cle57, Olv67, SAM08]. Lobatto [PDP12b]. locally [BCG74]. location [Row69]. Logarithmic [CLOT86, Dom14, Dun73b]. London [Cle63, W.63]. Look [Cle78]. Lord [JHC^{+}14, MC05b, MC05a]. lower [Tal76]. Ltd [Cle65]. Luke [Cle72].

Machinery [CO51]. machines [dGR95]. Maclaurin [Smi65]. Macmillan [Cle63]. Majesty [W.63]. manipulating [LM76]. MAPLE [GF92]. maps [HT08]. Markoff [Peh78]. Markov [Sch78]. Markov-type [Sch78]. Martin [Cle67]. Mathematical [Cle62a, Cle62b, W.63, CO55]. Mathematician [Olv04, Cle04, CMM16]. matrix [Bar98]. MDCT [NF03]. Mean [MP72]. means [Orl77b]. measures [CE09]. meilleur [MP72]. Memories [Cle92a]. memory [dGR95]. Method [Cle65, ES88, Oli77, Tua73, ATAAK17, Bra73, Bra92, CC60, De 75, Ell73, EL00, EL06, GC24, Hol69, HFWW22, Hun70, Imh63, KKY07, Loc69, Ma15, Oli71, Oli72, PB83, SAM08, SHN20, Smi65, Tay70, Ura69]. Methodology [Gen72b]. Methods [CGM^{+}61, Fil64, Cle67, Mas70, MV97, Oli75, Smi65, Sti65, XCWB11, XCWB13]. Mills [RP63]. minimax [Coh73]. Mixed [PDP12a, DD11, EL06, PDP12b]. modeling [LT92]. Modern [CGM^{+}61]. Modification [Håv69]. Modified [Bar13, Maj21a, Oli77, PB83, Pie87, ATAAK17, CL14, HW77]. Moment [RV07, VR11]. moments [JKH^{+}17]. moyenne [MP72]. MR2846756 [XCWB13]. Multi [Hay73]. Multi-dimensional [Hay73]. multiple [Lin76]. Multivariate [BCG74].

N.J [Cle67]. National [W.63]. near [Coh73]. nesting [Lin76]. Nets [Cle65]. Networks [GW22, GW23]. Neural [GW22, GW23]. Newbery [Oli80]. Newton [DD11, Tay70]. no [CMM16]. nodes [CS78]. nonlinear [CN63, Maj21a]. Nontensorial [SVZ08]. norm [Not16]. normal [OR78]. norms [Coh73]. Note [Cle55, BS71, Küt84]. Notes [BRCG55, CO51]. numeric [GF92]. Numerical [Ano78, CL23, CO51, CE60, Cle63, Cle68a, LPM81, PTVF92a, SAM08, Sut19c, Tur89, Ura69, XCWB13, Cle57, CC60, Cle60b, Cle67, Imh63, KEKT21, Mas70, Olv04, Pie87, SH11]. Nyström [CL14, KKR03].

O1 [Ged79]. Obituary [Cle92b]. obscure [CMM16]. obtaining [Mat72]. odd [Sma85]. off [Pry81]. Office [W.63]. one [Sal73]. ones [Orl77a]. Operations [CO87]. operators [CS78]. Optimal [Tay70, NTT15, SHN20]. Order [GCP67, CP66, Olv67, Pet94]. ordinary [BS71, Cle60b, CN63, Oli69, SAM08]. Orr [CE60]. Orthogonal [Mas70, Bar02, Hol69, Sal73]. oscillating [Ada87]. oscillator [Maj21a]. oscillatory [DGS11, DGK13, Dom14, GI17, GC24, KEKT21, LX19, Maj17, Maj19, Maj21a, XCWB11, XCWB13, XHC15]. Other [RP63, Han73, Küt84]. Oxford [Cle65, Cle68a].

p [W.63]. packaging [MP72]. Padé [Bak74, MC05b, MC05a, dB76]. pages [Cle63, Cle65, Cle67, Cle68a, Cle66a]. Papers [BRCG55]. Parabolic [Cle65]. Parallel [Tur89, dGR95]. parity [Boy23]. Parker [Cle68a]. Part [MC05b]. PDEs [NTT15]. performance [LRD19]. Pergamon [Cle65]. Permissable [Rou70]. Phase [Sti65]. phenomenon [WT07]. Physics [W.63]. Piecewise [BdBSW66, CL14, Pat76]. plane [CM83]. Point [CO84]. points [AN91, DGK13, Maj17, Oli75, Row69, SS78a, SS78b, SS80]. points-convergence [SS78a]. Pol [Cle66b]. Polynomial [Cle54, Cle64, Coh73, DK71, MC05b, RP63, And73, Bar02, Bar03, Dun75, FGR76, HW77, Has01, Mas70, MC05a, Neu74, New74, Oli80, Pat76, Sti65]. polynomial-rational [Dun75]. polynomiale [FGR76]. Polynomials [BdBSW66, Cle68a, Gau79, KR79, AGR93, CO55, Hay73, Hol69, Jai74, LM76, Luk76, MV97, Mat72, Mur71, Rou70, RT72, Row69, Sch78, ST69, Tal76, Tay68]. positive [Mat72]. Positivity [HST93]. Power [Gau79]. pp [Cle72]. practical [CMM16, Oli71]. preliminary [Rie71]. Prentice [Cle67]. Prentice-Hall [Cle67]. presence [Maj17]. Press [Cle65, Cle68a, Cle72]. Primitive [Fil64]. problem [TW74]. problems [Sch78, SHN20, Ura69]. procedure [Lin76]. Processing [Tur89]. Product [AN91, SS78a, SS78b, SS80, LPM81, Pet94]. Product-integration [SS78a, SS78b]. Professor [Cle92b]. Projections [Neu74]. properties [Bak74, Bra72, SS78a, SS78b, Smi15].

quadratique [MP72]. Quadrature [Bar13, Cha68, FW66, Ged79, Gen72a, Håv69, Kam71a, PDP12a, Smi65, Tre08, Wal06, XB13, Ada87, AF84, Boy23, BP75, Bra73, Bra92, CE09, CL14, CL23, DD11, Den86, EJJ08, EHM81, FLR93, GI17, Gen72b, Gen72c, Goo73, HT08, HTN83, HST93, HS15, HS05, Jay74, Kam71b, KKR03, KKY07, Kus72, LX19, Loc69, Ma15, Mar86, Not16, OS68, Oli71, Oli72, Pat76, PDP12b, Rie71, RJ71, RJ72, Smi82, SH11, Smi15, Sut19a, Sut19b, Wan18, WT07, XB12, XHW14, XHC15, XrL11]. quadratures [Küt84]. Quadraturverfahren [Bra73, Loc69]. quantities [Orl77b]. quantum [SAM08]. quasi [NTT15]. quasi-optimal [NTT15]. questions [TW74].

random [LRD19, NTT15]. ranges [RT72, ST69, Tay68, Tay69]. rapidity [CS78]. rate [Wan18]. Rates [XB13, XB12]. Ratio [RP63]. Rational [CL74, Cle74, WBR82, Dun75, XrL11]. rationals [TW74]. Rayleigh [BdBSW66]. Real [PDP12a, FGR76, PDP12b, SV73]. recipes [PTVF92a]. reciprocals [CO86]. reconstruction [JKH^{+}17]. Recurrence [CO51, PTVF92b, RV07, VR11, AZK95, JKH^{+}17, NF03, Sal73]. Recursive [VR11]. réel [FGR76]. reference [CMM16]. regular [CL23]. Regularity [XB13, XB12]. regulated [RR71]. Reinsch [Oli80]. related [AN91, Jai74, Not16, OR78, SS78a, SS78b]. Relations [CO51, PTVF92b]. relative [Han73]. Remark [Ged79, Goo73, Mar86]. Remarks [FW66]. Rémès [Car72a, Car72b]. report [Rie71]. restricted [RT72, ST69, Tay68, Tay69]. Review [Cle65, Cle66a, Cle67, Cle68a]. Reviews [Cle63, Cle72, W.63]. ring [Sch78]. Ritz [BdBSW66]. robust [AN91]. Root [CT89]. roots [CO86]. Round [Pry81]. Round-off [Pry81]. Rounding [Bar02, Bar03]. Rule [PDP12a, CE09, HS15, KKY07, LX19, PDP12b, XrL11]. Rules [PDP12a, Wal06, Ada87, AN91, CL23, DD11, DGS11, DGK13, Dom14, HST93, Maj19, Maj21b, Maj21a, MFA22, PDP12b, Pet94, Som13, SH22, XHW14, vWGB19].

Satzes [Peh78]. Saul’yev [Cle65]. Scheme [FW66, Oli80, Sal73]. School [Tur89]. science [Ful77]. scientific [PTVF92a]. second [CS78, Jai74, Olv67, XrL11]. second-order [Olv67]. semi [SH21, SH22]. semi-infinite [SH21, SH22]. separation [CJ74]. Series [Cle55, Cle62a, Cle64, ES88, GCP67, Oli77, W.63, Bar02, Bas73, BS71, Boy23, Cle57, Cle60b, CN63, CP66, Cle66b, CL74, Ell68, Hol69, Hun70, Lin76, Luk76, Oli69, TE72, Ura69]. sets [Bra72, Mat72]. several [Rei77]. Short [BRCG55]. simulation [LT92]. singular [CE09, CL23, KEKT21, Maj19]. singularities [DGK13, Dom14, Wan18]. sizes [Han73]. Smolyak [HNU14]. smooth [CL14, NR96]. Snyder [Cle67]. Software [Ful77]. Solution [CO51, Cle57, Cle60b, CN63, Cle66b, EHM81, Oli69, SAM08, Ura69]. solutions [Olv67]. solving [Orl77b]. some [DD11]. somme [MP72]. Sommerfeld [CE60]. sparse [NTT15]. Special [CMW63, Cle72, Cle74, Deu76]. spectral [EL00, EL06, HFWW22]. speed [SV73]. spherical [LPM81]. spline [CN78]. splines [JS69, Sma85]. Sprague [De 75]. square [CO86, GC24, MP72]. squares [KKY07]. squaring [CT89]. Stability [DGS11, Bar98, MFA22, Smo02]. Stable [KR79, Maj19, Oli75, XHW14]. Stammfunktion [Fil64]. starting [Tay70]. stationary [DGK13, Maj17]. Stationery [W.63]. statistics [vWGB19]. strategy [Oli71]. Stroud [Cle63]. Study [Car72b]. Suboptimal [Sch79]. sum [MP72]. Summation [Cle55, Boy23, Deu76]. Summer [Tur89]. summing [Ell68]. sums [Dun75, Rei77]. Surface [CJ74, CH65, Cle68b]. Survey [COT89]. Symbolic [Wan92, GF92]. symbolic-numeric [GF92]. symmetric [CT88]. Symmetrical [LT92]. Symposium [Wan92]. System [Wan92, CT88]. systems [Orl77b]. Szego [AGR93].

Table [GS48, dB76]. Tables

[Cle62b, Cle72, W.63, CO55]. Talbot

[dGR95]. Tchebichef [Jai74]. Tchebychev [CS78]. tears [Pry81]. Technical [BRCG55]. technique [Oli69]. term [SH11]. terms [Han73]. Their [Cle72, RT72]. Theorem [Peh78]. theorems [And73]. Theory [Sut19c, Ful77]. Time [Has19, KR79]. topological [Bra72]. tout [FGR76]. tractability [HNU14]. transcendent [Orl77b]. transfer [DK71]. Transform [RV07, VR11, AZK95, HF11, KEKT21]. transformation [Gen72c, Oli80]. transformations [Maj19]. transformed [Ma15]. transforming [Orl77a]. transforms [Pie87, XCWB11, XCWB13]. translated [Cle63]. Treatment [CE60]. Truncated [Cle64, Boy23, SH21]. Tschebyscheff [Fil64]. Tschebyscheff-Approximation [Fil64]. Type [Cle65, MC05b, AF84, CS78, DD11, EL06, FLR93, HS05, LX19, Luk76, MV97, MC05a, Sch78, SH11, Smi15, SH22, XCWB11, XCWB13, XHW14, vWGB19]. ultraspherical [Han73]. unbounded [HW77]. unified [Bar03]. Uniform [CS78, Ell73, Neu74, Tal76]. unisolvent [RR71]. univariate [BCG74]. University [Cle68a]. unknown [Cle48, Cle80]. Unrestricted [CO86, CO80]. use [CO55]. Using [Boy23, VR11, AZK95, BS71, CL23, CT89, EHM81, JHC^{+}14, NF03, RV07].

V [Cle63, Cle65, W.63]. value [CMM16, Ura69]. variables [Rei77]. varisolvent [Bra72]. Verfahrens [Fil64]. via [HF11, JKH^{+}17]. vol [Cle72]. Volterra [EHM81, HFWW22].

W [LE72, W.63]. W. [Tod67]. w.r.t [CE09]. Walsh [Ric64]. weak [HNU14]. weight [HS05, SH11, Smi15, Som13]. weight-function [HS05]. weighting [Kus72]. weights [HST93]. which [CMM16]. whole [FGR76]. wide [JHC^{+}14]. Wiener [XrL11]. Wilkinson [Ano78, Cle04]. William [CLG^{+}03].

X [W.63]. xx [Cle72].

York [Cle63, Cle72]. Yudell [Cle72].

zeros [Boy23]. Zolotareff [Tod67]. zum [Bra73].

References

Adam:1987:AFE

[Ada87] Gh. Adam. Alternative formulation of the extended Clenshaw– Curtis quadrature rules for oscillating integrals. Revue Roumaine de Physique, 32(8):813–826, 1987. CODEN RRPQAN. ISSN 0035-4090.

Akrivis:1984:DQF

[AF84] G. Akrivis and K.-J. Förster. On the definiteness of quadrature formulae of Clenshaw–Curtis type. Computing, 33(3–4):363– 366, September 1984. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ammar:1993:ASP

[AGR93] Gregory S. Ammar, William B. Gragg, and Lothar Reichel. An analogue for Szegő polynomials of the Clenshaw algorithm. Journal of Computational and Applied Mathematics, 46(1–2): 211–216, June 14, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/037704279390296N.

Computational complex analysis.

Adam:1991:PIR

[AN91] Gh. Adam and A. Nobile. Product integration rules at Clenshaw–Curtis and related points: a robust implementation. IMA Journal of Numerical Analysis, 11(2):271–296, April 1991. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Andria:1973:CTI

[And73] George D. Andria. Convergence theorems for integral polynomial approximations. Journal of Approximation Theory, 7: 319–324, 1973. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Anonymous:1978:CDJ

[Ano78] Anonymous, editor. The Contribution of Dr. J. H. Wilkinson to Numerical Analysis, Symposium Proceedings Series No. 19. The Institute of Mathematics and its Applications, 1978.

Anonymous:1992:LF

[Ano92] Anonymous. Leslie Fox: 1918– 1992. Oxford University booklet., 1992. ii + 77 pp.

URL https://people.maths.
ox.ac.uk/wathen/fox/Leslie_
Fox_booklet_1994.pdf.  In-

cludes brief C.V. (completed 3 days before Leslie Fox’s death), and a bibliography of his publications. Memories of Leslie Fox by C. W. Clenshaw, pp. 7–10.

Al-Towaiq:2017:MAC

[ATAAK17] Mohammad Al-Towaiq, Marwan Alquran, and Osama Al-Khazaleh. A modified algorithm for the Clenshaw–Curtis method. Journal of Information & Optimization Sciences, 38(3–4):455– 469, 2017. CODEN JIOSDC. ISSN 0252-2667.

Aburdene:1995:CDC

[AZK95] M. F. Aburdene, Jianqing Zheng, and R. J. Kozick. Computation of discrete cosine transform using Clenshaw’s recurrence formula. IEEE Signal Processing Letters, 2(8):155–156, August 1995. CO-DEN ISPLEM. ISSN 1070-9908 (print), 1558-2361 (electronic).

Baker:1974:CIC

[Bak74] George A. Baker, Jr. Certain invariance and convergence properties of the Padé approximant. Rocky Mountain Journal of Mathematics, 4(2):141– 149, June 1974. CODEN RMJ-MAE. ISSN 0035-7596 (print), 1945-3795 (electronic).

Barrio:1998:MAS

[Bar98] Roberto Barrio. A matrix analysis of the stability of the Clenshaw algorithm. Extracta Mathematicae, 13(1):21–26, 1998. ISSN

0213-8743. URL http://eudml.
org/doc/38560.

Barrio:2002:REB

[Bar02] Roberto Barrio. Rounding error bounds for the Clenshaw and Forsythe algorithms for the evaluation of orthogonal polynomial series. Journal of Computational and Applied Mathematics, 138 (2):185–204, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704270100382X.

Barrio:2003:URE

[Bar03]  Roberto Barrio.  A unified

rounding error bound for polynomial evaluation. Advances in Computational Mathematics, 19(4):385–399, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:/

/link.springer.com/article/
10.1023/A:1024203520270.

Barden:2013:MCC

[Bar13] Jeffrey Michael Barden. A modified Clenshaw–Curtis quadrature algorithm. M.S. in applied mathematics, Worcester Polytechnic Institute, Worcester, MA, USA, April 24, 2013. iv + 29 pp.

Basu:1973:DCS

[Bas73] N. K. Basu. On double Chebyshev series approximation. SIAM Journal on Numerical Analysis, 10(3):496–505, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Birkhoff:1974:MAL

[BCG74] Garrett Birkhoff, James C.

Cavendish, and William J. Gordon. Multivariate approximation by locally blended univariate interpolants. Proc. Natl. Acad. Sci. USA, 71(9):3423–3425, September 1, 1974. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Birkhoff:1966:RRA

[BdBSW66] Garrett Birkhoff, C. de Boor, B. Swartz, and B. Wendroff. Rayleigh–Ritz approximation by piecewise cubic polynomials. SIAM Journal on Numerical Analysis, 3(2):188–203, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boyd:2023:UPA

[Boy23] John P. Boyd. Using parity to accelerate Hermite function computations: zeros of truncated Hermite series, Gaussian quadrature and Clenshaw summation. Mathematics and Computers in Simulation, 207:521–532, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Branders:1975:ECC

[BP75] Maria Branders and Robert Piessens. An extension of Clenshaw–Curtis quadrature. Journal of Computational and Applied Mathematics, 1(1): 55–65, March 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X75900091.

Braess:1972:TPS

[Bra72] D. Braess. On topological properties of sets admitting varisolvent functions. Proceedings of the American Mathematical Society, 34(2):453–456, August 1972. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Brass:1973:FQC

[Bra73] Helmut Braß. Eine Fehlerabschätzung zum Quadraturverfahren von Clenshaw und Curtis. (German) [An error estimation of the quadrature method of Clenshaw and Curtis]. Numerische Mathematik, 21(5):397–403, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brass:1992:EEC

[Bra92] Helmut Braß. Error estimation for the Clenshaw–Curtis quadrature method. Abhandlungen der Braunschweigischen Wissenschaftlichen Gesellschaft, 43:45–53, 1992. CODEN AB-WGAZ. ISSN 0068-0737.

Brenner:1955:TNS

[BRCG55] J. L. Brenner, G. W. Reitwiesner, C. W. Clenshaw, and I. J. Good. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 9 (51):117–121, July 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Boggs:1971:NIO

[BS71] R. A. C. Boggs and F. J. Smith.

A note on the integration of ordinary differential equations using Chebyshev series. The Computer Journal, 14(3):270–271, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/
content/14/3/270.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_03/140270.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/270.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/271.tif.

Carasso:1972:DHA

[Car72a] C. Carasso. Densité des hypothèses assurant la convergence de l’algorithme de Rémès. (French) [Density of hypotheses ensuring the convergence of the Rémès algorithm]. Revue Française d’Automatique, Informatique et Recherche Opérationnelle, 6:69– 84, 1972. ISSN 0035-3035.

Carasso:1972:EAR

[Car72b] Claude Carasso. Étude de l’algorithme de Rémès en l’absence de condition de Haar (French) [study of the algorithm Rémès in the absence of Haar condition]. Numerische Mathematik, 20(2):165–178, April 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clenshaw:1960:MNI

[CC60] C. W. Clenshaw and A. R. Curtis. A method for numerical integration on an automatic computer. Numerische Mathematik, 2(1):197–205, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clenshaw:1960:NTO

[CE60] C. W. Clenshaw and D. Elliott. A numerical treatment of the Orr–Sommerfeld equation in the case of a laminar jet. Quarterly Journal of Mechanics and Applied Mathematics, 13(3):300– 313, 1960. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Calabro:2009:ECC

[CE09] F. Calabrò and A. Corbo Esposito. An evaluation of Clenshaw– Curtis quadrature rule for integration w.r.t. singular measures. Journal of Computational and Applied Mathematics, 229 (1):120–128, July 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042708005256.

Clenshaw:1961:MCM

[CGM^{+}61] C. W. Clenshaw, E. T. Goodwin, D. W. Martin, G. F. Miller, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of National Physical Laboratory. Notes on Applied Science. Her Majesty’s Stationery Office, London, UK, second edition, 1961. vi + 170 pp. LCCN QA297 .T4 1961.

Clenshaw:1965:CSF

[CH65] C. W. Clenshaw and J. G. Hayes.

Curve and surface fitting. Journal of the Institute of Mathematics and its Applications, 1(2): 164–183, June 1965. CODEN JMTAA8. ISSN 0020-2932.

Chawla:1968:EEC

[Cha68] M. M. Chawla. Error estimates for the Clenshaw–Curtis quadrature. Mathematics of Computation, 22(103):651–656, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Call:1974:SFS

[CJ74] E. S. Call and F. F. Judd. Surface fitting by separation. Journal of Approximation Theory, 12: 283–290, 1974. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Clenshaw:1974:RAC

[CL74] C. W. Clenshaw and K. Lord. Rational approximations from Chebyshev series. In Studies in Numerical Analysis (papers in Honour of Cornelius Lanczos on the Occasion of His 80th Birthday), pages 95–113. Royal Irish Academy, Dublin, Ireland, 1974.

Chen:2014:MNC

[CL14] Qiong-Sheng Chen and Fu-Rong Lin. A modified Nyström– Clenshaw–Curtis quadrature for integral equations with piecewise smooth kernels. Applied Numerical Mathematics: Transactions of IMACS, 85:77–89, November 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2023:NCR

[CL23] Linchong Chen and Xiaolin Li. Numerical calculation of regular and singular integrals in boundary integral equations using Clenshaw–Curtis quadrature rules. Engineering Analysis with Boundary Elements, 155:25–37, 2023. ISSN 0955-7997 (print), 1873-197X (electronic).

Clenshaw:1948:U

[Cle48] Charles William Clenshaw. [unknown]. Ph.D. dissertation, University of London, London, UK, 1948.

Clenshaw:1954:PAE

[Cle54] C. W. Clenshaw. Polynomial approximations to elementary functions. Mathematical Tables and Other Aids to Computation, 8 (47):143–147, July 1954. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1955:NSC

[Cle55] C. W. Clenshaw. A note on the summation of Chebyshev series. Mathematical Tables and Other Aids to Computation, 9 (51):118–120, July 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1957:NSL

[Cle57] C. W. Clenshaw. The numerical solution of linear differential equations in Chebyshev series. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 53(1): 134–149, January 1957. CODEN PCPSA4. ISSN 0008-1981.

Clenshaw:1960:CFD

[Cle60a] C. W. Clenshaw. Curve fitting with a digital computer. The Computer Journal, 2(4): 170–173, January 1960. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/

2/4/170.full.pdf+html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/020170.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_
04/tiff/170.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_04/tiff/
171.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/172.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
02/Issue_04/tiff/173.tif.

Clenshaw:1960:NSO

[Cle60b] C. W. Clenshaw. The numerical solution of ordinary differential equations in Chebyshev series. In Symposium on the Numerical Treatment of Ordinary Differential Equations, Integral and Integro-differential Equations (Rome, 1960), pages 222–227. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1960.

Clenshaw:1962:CSM

[Cle62a] C. W. Clenshaw. Chebyshev Series for Mathematical Functions, volume 5 of National Physical Laboratory. Mathematical tables. Her Majesty’s Stationery Office, London, UK, 1962. iv + 36 pp. Department of Scientific and Industrial Research.

Clenshaw:1962:MT

[Cle62b] C. W. Clenshaw. Mathematical Tables. National Physical Laboratory. Mathematical tables. Her Majesty’s Stationery Office, London, UK, 1962. 36 pp.

Clenshaw:1963:BRN

[Cle63] C. W. Clenshaw. Book reviews: Numerical analysis: Approximate Calculation of Integrals, by V. I. Krylov, translated by Arthur H. Stroud, 1962; 357 pages. (New York: The Macmillan Company, $7.50. London: The Macmillan Company, New York). The Computer Journal, 6 (2):207–208, August 1963. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
6/2/207.full.pdf+html.

Clenshaw:1964:CBP

[Cle64] C. W. Clenshaw. A comparison of “Best” polynomial approximations with truncated Chebyshev series expansions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(1):26–37, January 1964. ISSN 0887-459X (print), 2168-3581 (electronic).

Clenshaw:1965:BRI

[Cle65] C. W. Clenshaw. Book review: Integration of Equations of Parabolic Type by the Method of Nets, V. K. Saul’yev, 1964; 346 pages. (Oxford: Pergamon Press Ltd.). The Computer Journal, 8(2):156, February 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.

                                        oxfordjournals.org/content/
                                        8/2/156.full.pdf+html.

Clenshaw:1966:BRA

[Cle66a] C. W. Clenshaw. Book review: Approximation of Functions, edited by H. Garabedian, 1965; 215 pages. The Computer Journal, 9(3):285, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.
                                        oxfordjournals.org/content/
                                        9/3/285.full.pdf+html.

Clenshaw:1966:SVP

[Cle66b] C. W. Clenshaw. The solution of van der Pol’s equation in Chebyshev series. In Numerical Solutions of Nonlinear Differential Equations (Proceedings of the Advanced Symposium, Madison, Wisconsin, 1966), pages 55– 63. Wiley, New York, NY, USA, 1966.

Clenshaw:1967:CMN

[Cle67] C. W. Clenshaw. Book review: Chebyshev methods in numerical approximation: by Martin A. Snyder. 114 pages, 6 × 9 in. Englewood Cliffs, N.J., Prentice-Hall, Inc., 1966. Journal of The Franklin Institute, 284(5):338–??, November 1967. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic). URL https:

                                        //api.semanticscholar.org/
                                        CorpusID:117920278.

Clenshaw:1968:BRC

[Cle68a] C. W. Clenshaw. Book review: Chebyshev Polynomials in Numerical Analysis, by L. Fox and I. B. Parker, 1969; 205 pages (Oxford University Press). The Computer Journal, 11(3):310, March 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/
content/11/3/310.full.pdf+
html.

Clenshaw:1968:CSF

[Cle68b] C. W. Clenshaw. Curve and surface fitting. Report NPL Ma 72, National Physical Laboratory Division of Numerical Applied Mathematics, Teddington, UK, 1968. 21 pp.

Clenshaw:1972:RDT

[Cle72]  C. W. Clenshaw.  Reviews

and descriptions of tables and books: Yudell L. Luke, The Special Functions and Their Approximations, Academic Press, New York, 1969, vol. I, xx + 349 pp., vol. II, xx + 485 pp., 23 cm. Mathematics of Computation, 26(117):297– 299, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.jstor.org/
stable/2004747.

Clenshaw:1974:RAS

[Cle74] C. W. Clenshaw. Rational approximations for special functions. In Software for Numerical Mathematics (Proceedings of a Conference, Institute of Mathematics and Applications, Loughborough University Tech., Loughborough, 1973), pages 275–284, 301. Academic Press, New York, USA, 1974. With discussion by J. A. Grant and C. W. Clenshaw.

Clenshaw:1978:NLA

[Cle78] C. W. Clenshaw. A new look at error analysis. In Anonymous [Ano78], pages 54–75.

Clenshaw:1980:U

[Cle80] Charles William Clenshaw. [unknown]. D.Sc., University of London, London, UK, 1980.

Clenshaw:1992:MLF

[Cle92a]  C. W. Clenshaw.  Memo-

ries of Leslie Fox. In LF-1992 [Ano92], pages 8–10.

                                        URL https://people.maths.
                                        ox.ac.uk/wathen/fox/Leslie_
                                        Fox_booklet_1994.pdf.  In-

cludes brief C.V. (completed 3 days before Leslie Fox’s death), and a bibliography of his publications. Memories of Leslie Fox by C. W. Clenshaw, pp. 7–10.

Clenshaw:1992:OPL

[Cle92b] Charles Clenshaw. Obituary: Professor Leslie Fox. The Independent, ??(??):??, August 10, 1992. URL https://

                                        www.independent.co.uk/news/
                                        people/obituary-professor-
                                        leslie-fox-1539620.html.

Clenshaw:2004:WJH

[Cle04] C. W. Clenshaw. Wilkinson, James Hardy (1919–1986), mathematician. In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, September 2004.

Cheney:2003:WAL

[CLG^{+}03] E. Ward Cheney, Jeremey Levesley, Charles Goldie, Charles W. Clenshaw, Graham Jameson, Peter Turner, Rick Beatson, and Manfred v. Golitschek. William Allan Light: (19 April 1950– 8 December 2002). Journal of Approximation Theory, 123(1):1– 12, July 2003. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Clenshaw:1986:GEL

[CLOT86] C. W. Clenshaw, Daniel W.

Lozier, Frank W. J. Olver, and Peter R. Turner. Generalized exponential and logarithmic functions. Computers and Mathematics and Applications, 12 (Part 2)(5–6):1091–1101, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902336.

Coleman:1983:CEB

[CM83] J. P. Coleman and A. J. Monaghan. Chebyshev expansions for the Bessel function J_{n}(z) in the complex plane. Mathematics of Computation, 40(161): 343–366, January 1983. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Clenshaw:2016:OVR

[CMM16] C. W. Clenshaw, K. Miller, and W. Meecham. obscure the value of a reference book which no practical mathematician can afford. Unknown, 2016. URL https:

//www.semanticscholar.org/
paper/obscure-the-value-of-
a-reference-book-which-no-
can-Clenshaw-Miller/cd7c67a84fe815bd76ab7b48997e345918fae72d.

Clenshaw:1963:ASF

[CMW63] C. W. Clenshaw, G. F. Miller, and M. Woodger. Algorithms for special functions I. Numerische Mathematik, 4(1):403– 419, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). From the Web page: “Editor’s note. In this fascicle, prepublication of algorithms from the Special Function series of the Handbook for Automatic Computation is started. Algorithms are published in ALGOL 60 reference language [1] as approved recently by the IFIP.”.

Clenshaw:1963:SNO

[CN63] C. W. Clenshaw and H. J. Norton. The solution of nonlinear ordinary differential equations in Chebyshev series. The Computer Journal, 6(1):88–92, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
6/1/88.full.pdf+html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/060088.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/88.tif;   http://www3.
oup.co.uk/computer_journal/

hdb/Volume_06/Issue_01/tiff/
89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/90.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/91.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_
01/tiff/92.tif.

Clenshaw:1978:CSA

[CN78] C. W. Clenshaw and B. Negus.

The cubic X-spline and its application to interpolation. Journal of the Institute of Mathematics and its Applications, 22(1):109– 119, August 1978. CODEN JM-TAA8. ISSN 0020-2932.

Clenshaw:1951:NNA

[CO51] C. W. Clenshaw and F. W. J.

Olver. Notes on numerical analysis — 3: Solution of differential equations by recurrence relations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 5(33):34–39, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1955:UEP

[CO55] C. W. Clenshaw and Frank W. J. Olver. The use of economized polynomials in mathematical tables. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 51(4):614–628, October 1955. CODEN PCPSA4. ISSN 0008-1981.

Clenshaw:1980:UAE

[CO80] C. W. Clenshaw and Frank W. J. Olver. An unrestricted algorithm for the exponential function. SIAM Journal on Numerical Analysis, 17(2):310–331, April 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clenshaw:1984:BFP

[CO84] C. W. Clenshaw and Frank W. J. Olver. Beyond floating point. Journal of the ACM, 31(2):319– 328, March 30, 1984. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clenshaw:1986:UAR

[CO86] C. W. Clenshaw and Frank W. J. Olver. Unrestricted algorithms for reciprocals and square roots. BIT (Nordisk tidskrift for informationsbehandling), 26(4):476– 492, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Clenshaw:1987:LIA

[CO87] C. W. Clenshaw and Frank W. J. Olver. Level-index arithmetic operations. SIAM Journal on Numerical Analysis, 24(2):470–485, April 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cohen:1973:PAF

[Coh73] Edgar A. Cohen, Jr. Polynomial approximation of a function and its first derivative in near minimax norms. Mathematics of Computation, 27(124): 817–827, October 1973. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Clenshaw:1989:LIA

[COT89] C. W. Clenshaw, F. W. J. Olver, and Peter R. Turner. Level-index arithmetic: An introductory survey. In Turner [Tur89], pages 95– 168. CODEN LNMAA2. ISBN 3-540-51645-X. Lectures from the Third S.E.R.C. Numerical Analysis Summer School held at the University of Lancaster, Lancaster, 1987.

Clenshaw:1966:CSB

[CP66] C. W. Clenshaw and Susan M.

Picken. Chebyshev series for Bessel functions of fractional order, volume 8 of National Physical Laboratory. Mathematical tables. Her Majesty’s Stationery Office, London, UK, 1966. iii + 53 pp.

Chauhan:1978:UCR

[CS78] B. P. S. Chauhan and K. B. Srivastava. Uniform convergence and rapidity of convergence of Grunwald-type operators on extended Tchebychev nodes of second kind. Indian Journal of Pure and Applied Mathematics, 9(12): 1337–1343, 1978. ISSN 0019-5588 (print), 0975-7465 (electronic).

Clenshaw:1988:SLI

[CT88] C. W. Clenshaw and Peter R.

Turner. The symmetric levelindex system. IMA Journal of Numerical Analysis, 8(4):517– 526, October 1988. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Clenshaw:1989:RSU

[CT89] C. W. Clenshaw and Peter R. Turner. Root squaring using level-index arithmetic. Computing, 43(2):171–185, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deBruin:1976:CPT

[dB76]  M. G. de Bruin.  Conver-

gence in the Padé table for 1F_{1}(1; c; x). Indagationes Mathematicæ, 38(5):408–418, 1976. CODEN IMTHBJ. ISSN 0019-3577 (print), 1872-6100 (electronic). Nederl. Akad. Wetensch. Proc. Ser. A 79.

Dash:2011:ISC

[DD11] Rajani Ballav Dash and Debasish Das. Identification of some Clenshaw–Curtis quadrature rules as mixed quadrature of Fejer and Newton–Cote type of rules. International Journal of Mathematical Sciences & Applications, 1(3):1493–1496, 2011. ISSN 2230-9888.

DeKerf:1975:IMS

[De 75] Joseph L. F. De Kerf. The interpolation method of Sprague– Karup. Journal of Computational and Applied Mathematics, 1(2):101–110, June 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Denisenko:1986:CCQ

[Den86] P. N. Denisenko. The Clenshaw and Curtis quadrature formula. Izvestiia Vysshikh Uchebnykh Zavedenii. Matematika, 4: 71–73, 85, 1986. CODEN IVUMBY. ISSN 0021-3446 (print), 2076-4626 (electronic).

Deuflhard:1976:ASC

[Deu76] P. Deuflhard. On algorithms for the summation of certain special functions. Computing, 17 (1):37–48, March 1976. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dominguez:2013:FCC

[DGK13] V. Domínguez, I. G. Graham, and T. Kim. Filon–Clenshaw– Curtis rules for highly oscillatory integrals with algebraic singularities and stationary points. SIAM Journal on Numerical Analysis, 51(3):1542–1566, January 2013. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deRosa:1995:PTA

[dGR95] M. A. de Rosa, G. Giunta, and M. Rizzardi. Parallel Talbot’s algorithm for distributed memory machines. Parallel Computing, 21(5):783–80, May 10, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1995&volume=21&issue=
5&aid=974.

Dominguez:2011:SEE

[DGS11] V. Domínguez, I. G. Graham, and V. P. Smyshlyaev. Stability and error estimates for Filon– Clenshaw–Curtis rules for highly oscillatory integrals. IMA Journal of Numerical Analysis, 31 (4):1253–1280, April 2011. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Dierieck:1971:PCA

[DK71] C. Dierieck and Y. Kamp. Polynomial Chebyshev approximation of a complex transfer function. Computing, 8(3–4):335– 342, September 1971. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dominguez:2014:FCC

[Dom14] Víctor Domínguez. Filon– Clenshaw–Curtis rules for a class of highly-oscillatory integrals with logarithmic singularities. Journal of Computational and Applied Mathematics, 261:299–319, May 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Dunham:1973:CA

[Dun73a] Ch. B. Dunham. Chebyshev approximation by A^{∗}Φ(Bx). Zeitschrift für Angewandte Mathematik und Mechanik, 53(6):353– 354, 1973. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Dunham:1973:CAL

[Dun73b] Ch. B. Dunham. Chebyshev approximation by logarithmic families. Zeitschrift für Angewandte Mathematik und Mechanik, 53 (6):352–353, 1973. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Dunham:1975:CAP

[Dun75] Charles B. Dunham. Chebyshev approximation by polynomialrational sums. Zeitschrift für Angewandte Mathematik und Mechanik, 55(12):758–759, 1975. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Evans:1981:ISV

[EHM81] G. A. Evans, J. Hyslop, and A. P. G. Morgan. Iterative solution of Volterra integral equations using Clenshaw– Curtis quadrature. Journal of Computational Physics, 40 (1):64–76, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0021999181901996.

Elliott:2008:CCG

[EJJ08] David Elliott, Barbara M. Johnston, and Peter R. Johnston. Clenshaw–Curtis and Gauss– Legendre quadrature for certain boundary element integrals. SIAM Journal on Scientific Computing, 31(1):510–530, January 2008. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Elnagar:2000:CCS

[EL00] Gamal N. Elnagar and Ken S. Li. Clenshaw–Chebyshev spectral method for Hammerstein equations. International Journal of Differential Equations and Applications, 1(3):307–321, 2000. ISSN 1311-2872.

Elnagar:2006:CCS

[EL06] Gamal N. Elnagar and Ken S. Li. Clenshaw–Chebyshev spectral method for mixed type Hammerstein equations. Far East Journal of Applied Mathematics, 23(2):135–145, 2006. ISSN 0972-0960.

Elliott:1968:EAA

[Ell68] David Elliott. Error analysis of an algorithm for summing certain finite series. Journal of the Australian Mathematical Society, 8(2):213–221, May 1968. CO-DEN JAUMAX. ISSN 0004-9735 (print), 2059-9234 (electronic).

Elliott:1973:DMA

[Ell73] D. Elliott. A direct method for “almost” best uniform approximations. In Error, approximation, and accuracy (Proceedings of a Seminar, Australian National University, Canberra, 1972), pages 129–143. University of Queensland Press, St. Lucia, Brisbane, QLD, Australia, 1973.

Ellacott:1988:CMG

[ES88] S. W. Ellacott and E. B. Saff. On Clenshaw’s method and a generalisation to Faber series. Numerische Mathematik, 52(5):

499–509, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freud:1976:APT

[FGR76] G. Freud, A. Giroux, and Q. I.

Rahman. Sur l’approximation polynomiale sur tout l’axe réel, (French) [On the polynomial approximation on the whole real axis]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(5): 961–967, October 1, 1976. CO-DEN CJMAAB. ISSN 0008-414X.

Filippi:1964:ATA

[Fil64] Siegfried Filippi. Angenäherte Tschebyscheff-Approximation einer Stammfunktion des Verfahrens von Clenshaw und Curtis. (German) [Chebyshev approximation of a primitive function of the Clenshaw–Curtis methods]. Numerische Mathematik, 6(1):320– 328, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Favati:1993:BEF

[FLR93] Paola Favati, Grazia Lotti, and Francesco Romani. Bounds on the error of Fejér and Clenshaw– Curtis type quadrature for analytic functions. Applied Mathematics Letters, 6(6):3–8, November 1993. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Fullerton:1977:SAA

[Ful77] L. Wayne Fullerton. Software for approximations or approximation theory as an experimental science. In Padé and rational approximation (Proceedings of the International Symposium, University of South Florida, Tampa, Florida, 1976), pages 399–411. Academic Press, New York, USA, 1977. ISBN 0-12-614150-9.

Fraser:1966:RCC

[FW66] W. Fraser and M. W. Wilson. Remarks on the Clenshaw–Curtis quadrature scheme. SIAM Review, 8(3):322–327, July 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautschi:1979:CPP

[Gau79] Walter Gautschi. The condition of polynomials in power form. Mathematics of Computation, 33(145):343–352, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gao:2024:BFC

[GC24] Jing Gao and Gaoqin Chang. A bivariate Filon–Clenshaw–Curtis method of the highly oscillatory integrals on a square. Journal of Computational and Applied Mathematics, 439:Paper No. 115599, March 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Goldstein:1967:CSB

[GCP67] Max Goldstein, C. W. Clenshaw, and Susan M. Picken. Chebyshev series for Bessel functions of fractional order. Mathematics of Computation, 21(99):509– ??, July 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Geddes:1979:RCC

[Ged79] K. O. Geddes. Remark on “Algorithm 424: Clenshaw–Curtis quadrature [O1]”. ACM Transactions on Mathematical Software, 5(2):240, June 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gen72a].

Gentleman:1972:AAC

[Gen72a] W. Morven Gentleman. ACM Algorithm 424: Clenshaw–Curtis quadrature [D1]. Communications of the ACM, 15(5):353–355, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Goo73, Ged79].

Gentleman:1972:ICCa

[Gen72b] W. Morven Gentleman. Implementing Clenshaw–Curtis quadrature. I. Methodology and experience. Communications of the ACM, 15(5):337–342, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gentleman:1972:ICCb

[Gen72c] W. Morven Gentleman. Implementing Clenshaw–Curtis quadrature. II. Computing the cosine transformation. Communications of the ACM, 15(5):343–346, May 1972. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Geddes:1992:HSI

[GF92] K. O. Geddes and G. J. Fee. Hybrid symbolic-numeric integration in MAPLE. In Wang [Wan92], pages 36–41. ISBN 0-89791-489-9. LCCN QA 76.95 I59 1992.

Gao:2017:GFC

[GI17] Jing Gao and Arieh Iserles. A generalization of Filon– Clenshaw–Curtis quadrature for highly oscillatory integrals. BIT (Nordisk tidskrift for informationsbehandling), 57(4):943–961, September 2017. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Good:1973:RAC

[Goo73] Albert J. Good. Remark on Algorithm 424: Clenshaw–Curtis quadrature [D1]. Communications of the ACM, 16(8):490, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gen72a].

Goodwin:1948:T

[GS48] E. T. Goodwin and J. Staton. Table of du. Quarterly Journal of Mechanics and Applied Mathematics, 1(1):319–326, January 1948. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Guo:2022:CGN

[GW22] Yuhe Guo and Zhewei Wei. Clenshaw graph neural networks. Technical report, Renmin University of China, Beijing, China, October 2022. 1–

10 pp. URL https://www.
researchgate.net/publication/
364953685_Clenshaw_Graph_Neural_

Networks. Published in [GW23].

Guo:2023:CGN

[GW23] Yuhe Guo and Zhewei Wei. Clenshaw graph neural networks. In Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. ACM Press, New York, NY 10036, USA, August 2023.

Handscomb:1973:RST

[Han73] D. C. Handscomb. The relative sizes of the terms in Chebyshev and other ultraspherical expansions. Journal of the Institute of Mathematics and its Applications, 11(2):241–246, April 1973. CODEN JMTAA8. ISSN 0020-2932.

Hasegawa:2001:EAC

[Has01] Takemitsu Hasegawa. Error analysis of Clenshaw’s algorithm for evaluating derivatives of a polynomial. BIT Numerical Mathematics, 41(5):1019–1028, December 2001. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=41&issue=5&spage=1019.

BIT 40th Anniversary Meeting.

Hashemi:2019:ECE

[Has19]  Behnam Hashemi.  Enclos-

ing Chebyshev expansions in linear time. ACM Transactions on Mathematical Software, 45(3):27:1–27:33, July 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3319395.

Haavie:1969:MCC

[Håv69] T. Håvie. On a modification of the Clenshaw–Curtis quadrature formula. BIT (Nordisk tidskrift for informationsbehandling), 9(4):338–350, December 1969. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0006-3835&
                                        volume=9&issue=4&spage=338.

Hays:1973:MDE

[Hay73]  Richard O. Hays.  Multi-

dimensional extension of the Chebyshev polynomials. Mathematics of Computation, 27(123): 621–624, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hashemiparast:2011:ALT

[HF11] S. M. Hashemiparast and H. Fallahgoul. Approximation of Laplace transform of fractional derivatives via Clenshaw–Curtis integration. International Journal of Computer Mathematics, 88 (6):1224–1238, 2011. CODEN IJCMAT. ISSN 0020-7160.

Huang:2022:CCS

[HFWW22] Chaolan Huang, Chunhua Fang, Jianyu Wang, and Zhengsu Wan. On Clenshaw–Curtis spectral collocation method for Volterra integral equations. Journal of Applied Mathematics and Informatics, 40(5–6):983–993, 2022. ISSN 2734-1194 (print), 2234-8417 (electronic).

Hinrichs:2014:WTC

[HNU14] Aicke Hinrichs, Erich Novak, and Mario Ullrich. On weak tractability of the Clenshaw– Curtis Smolyak algorithm. Journal of Approximation Theory, 183:31–44, July 2014. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Holdeman:1969:MAF

[Hol69] Jonas T. Holdeman, Jr. A method for the approximation of functions defined by formal series expansions in orthogonal polynomials. Mathematics of Computation, 23(106):275– 287, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunter:2005:QFC

[HS05] D. B. Hunter and H. V. Smith. A quadrature formula of Clenshaw– Curtis type for the Gegenbauer weight-function. Journal of Computational and Applied Mathematics, 177(2):389–400, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042704004352.

Hasegawa:2015:EEC

[HS15] Takemitsu Hasegawa and Hiroshi Sugiura. Error estimate for a corrected Clenshaw–Curtis quadrature rule. Numerische Mathematik, 130(1):135–149, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

                                        /link.springer.com/article/
                                        10.1007/s00211-014-0660-y.

Hasegawa:1993:PWE

[HST93] Takemitsu Hasegawa, Hirosi Sugiura, and Tatsuo Torii. Positivity of the weights of extended Clenshaw–Curtis quadrature rules. Mathematics of Computation, 60(202):719–734, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hale:2008:NQF

[HT08] Nicholas Hale and Lloyd N. Trefethen. New quadrature formulas from conformal maps. SIAM Journal on Numerical Analysis, 46(2):930–948, January 2008. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

                                        tronic).  URL ttps://www.
                                        jstor.org/stable/40233238.

Hasegawa:1983:GCI

[HTN83] Takemitsu Hasegawa, Tatsuo Torii, and Ichizo Ninomiya. Generalized Chebyshev interpolation and its application to automatic quadrature. Mathematics of Computation, 41(164):537–553, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunter:1970:CME

[Hun70]  D. B. Hunter.  Clenshaw’s

method for evaluating certain finite series. The Computer Journal, 13(4):378–381, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Habib:1977:DAF

[HW77] Anwar Habib and Abdul Wafi.

Degree of approximation of functions by modified Bernstein polynomial on an unbounded interval. Indian Journal of Pure and Applied Mathematics, 8(6):691–695, 1977. ISSN 0019-5588 (print), 0975-7465 (electronic).

Imhof:1963:MNI

[Imh63] J. P. Imhof. On the method for numerical integration of Clenshaw and Curtis. Numerische Mathematik, 5(1):138– 141, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jaiswal:1974:PRT

[Jai74] D. V. Jaiswal. On polynomials related to Tchebichef polynomials of the second kind. Fibonacci Quarterly, 12:263–265, 1974. CODEN FIBQAU. ISSN 0015-0517.

Jayarajan:1974:EEG

[Jay74]  N. Jayarajan.  Error esti-

mates for Gauss–Chebyshev and Clenshaw–Curtis quadrature formulas. Calcolo: a quarterly on numerical analysis and theory of computation, 11(2):289–296, June 1974. CODEN CALOBK. ISSN 0008-0624 (print), 1126-5434 (electronic).

Jeong:2014:FAW

[JHC^{+}14] Yi-Ru Jeong, Ic-Pyo Hong, Heoung-Jae Chun, Yong Bae Park, Youn-Jae Kim, and Jong-Gwan Yook. Fast analysis over a wide band using Chebyshev approximation with Clenshaw– Lord approximation. In The 8th European Conference on Antennas and Propagation (EuCAP 2014). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 2014.

Jahid:2017:IMR

[JKH^{+}17] Tarik Jahid, Hicham Karmouni, Abdeslam Hmimid, Mhamed Sayyouri, and Hassan Qjidaa. Image moments and reconstruction by Krawtchouk via Clenshaw’s recurrence formula. In 2017 International Conference on Electrical and Information Technologies (ICEIT). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, November 2017.

Jerome:1969:S

[JS69] J. W. Jerome and L. L. Schumaker. On Lg-splines. Journal of Approximation Theory, 2: 29–49, 1969. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Kambo:1971:EBC

[Kam71a] N. S. Kambo. Error bounds for the Clenshaw–Curtis quadrature formulas. BIT (Nordisk tidskrift for informationsbehandling), 11(3):299–309, September 1971. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=11&issue=3&spage=299.

Kambo:1971:ECG

[Kam71b] N. S. Kambo. Error of certain Gauss quadrature formulas. Journal of the Institute of Mathematics and its Applications, 7 (3):303–307, June 1971. CODEN JMTAA8. ISSN 0020-2932.

Kayijuka:2021:CCA

[KEKT21] I. Kayijuka, Ş. M. Ege, A. Konuralp, and F. S. Topal. Clenshaw– Curtis algorithms for an efficient numerical approximation of singular and highly oscillatory Fourier transform integrals. Journal of Computational and Applied Mathematics, 385:Paper No. 113201, March 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Kang:2003:NCC

[KKR03] Sheon-Young Kang, Israel Koltracht, and George Rawitscher. Nyström–Clenshaw–Curtis quadrature for integral equations with discontinuous kernels. Mathematics of Computation, 72(242): 729–756, March 2003. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kim:2007:GCC

[KKY07] Changho Kim, Sang Dong Kim, and Jungho Yoon. Generalized Clenshaw–Curtis quadrature rule with application to a collocation least-squares method. Applied Mathematics and Computation, 190(1):781–789, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300307001129.

Kusterer:1979:SEP

[KR79] Roland Kusterer and Manfred Reimer. Stable evaluation of polynomials in time log n. Mathematics of Computation, 33(147): 1019–1031, July 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kussmaul:1972:CCQ

[Kus72] R. Kussmaul. Clenshaw–Curtis quadrature with a weighting function. Computing, 9(2): 159–164, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kutz:1984:NEE

[Küt84] Martin Kütz. A note on error estimates for the Clenshaw–Curtis and other interpolatory quadratures. Analysis, 4(1–2):45–51, January 1984. ISSN 0174-4747 (print), 2196-6753 (electronic).

Lam:1972:CCW

[LE72] Binh Lam and David Elliott.

On a conjecture of C. W. Clenshaw. SIAM Journal on Numerical Analysis, 9(1):44–52, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1976:GNP

[Lin76] E. J. Lin. Generalized nesting procedure of multiple series. Tamkang Journal of Mathematics, 7(2):195–199, 1976. ISSN 0049-2930 (print), 2073-9826 (electronic).

Lal:1976:CLM

[LM76] M. Lal and D. Moffatt. CHEBY-LAN — a language for manipulating Chebyshev polynomials. In Proceedings of the Fifth Manitoba Conference on Numerical Mathematics (University of Manitoba, Winnipeg, Manitoba, 1975), volume XVI of Congress. Numer., pages 439–448. Utilitas Math., Winnipeg, MB, Canada, 1976.

Locher:1969:FQC

[Loc69] F. Locher. Fehlerabschätzungen für das Quadraturverfahren von Clenshaw und Curtis. (German) [Error estimates for the quadrature method of Clenshaw and Curtis]. Computing, 4(4):304– 315, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lehman:1981:NEI

[LPM81] D. R. Lehman, William C. Parke, and L. C. Maximon. Numerical evaluation of integrals containing a spherical Bessel function by product integration. Journal of Mathematical Physics, 22 (7):1399–1413, July 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/
v22/i7/p1399_s1.

Loukrezis:2019:APL

[LRD19] Dimitrios Loukrezis, Ulrich Römer, and Herbert De Gersem. Assessing the performance of Leja and Clenshaw–Curtis collocation for computational electromagnetics with random input data. International Journal for Uncertainty Quantification, 9(1): 33–57, 2019. ISSN 2152-5080 (print), 2152-5099 (electronic).

Lozier:1992:SLI

[LT92] D. W. Lozier and P. R. Turner. Symmetrical level index arithmetic in simulation and modeling. Journal of research of the National Institute of Standards and Technology, 97(4):471– 485, July/August 1992. CO-DEN JRITEF. ISSN 1044-677X (print), 2165-7254 (electronic).

Luke:1976:EET

[Luk76] Yudell L. Luke. On the expansion of exponential type integrals in series of Chebyshev polynomials. In Theory of approximation, with applications (Proceedings of the Conference, University of Calgary, Calgary, Alberta, 1975; dedicated to the memory of Eckard Schmidt), pages 180– 199. Academic Press, New York, USA, 1976.

Liu:2019:CCT

[LX19] Guidong Liu and Shuhuang Xiang. Clenshaw–Curtis-type quadrature rule for hypersingular integrals with highly oscillatory kernels. Applied Mathematics and Computation, 340: 251–267, January 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2015:ICI

[Ma15]  Junjie Ma.  Implementing

the complex integral method with the transformed Clenshaw– Curtis quadrature. Applied Mathematics and Computation, 250(??):792–797, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0096300314013320.

Majidian:2017:FCC

[Maj17]  Hassan Majidian.  Filon–

Clenshaw–Curtis formulas for highly oscillatory integrals in the presence of stationary points.

Applied Numerical Mathematics: Transactions of IMACS, 117:87– 102, July 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Majidian:2019:SAF

[Maj19] Hassan Majidian. Stable application of Filon–Clenshaw–Curtis rules to singular oscillatory integrals by exponential transformations. BIT (Nordisk tidskrift for informationsbehandling), 59(1): 155–181, October 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Majidian:2021:MFC

[Maj21a] Hassan Majidian. Modified Filon–Clenshaw–Curtis rules for oscillatory integrals with a nonlinear oscillator. Electronic Transactions on Numerical Analysis, 54:276–295, 2021. ISSN 1068-9613 (print), 1097-4067 (electronic).

Majidian:2021:AEF

[Maj21b] Hassan Majidian. On adaptive extended Filon–Clenshaw–Curtis rules. In 2021 52nd Annual Iranian Mathematics Conference (AIMC). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 2021.

Marshak:1986:RCC

[Mar86] A. Marshak. A remark on the Clenshaw–Curtis quadrature formula. Eesti NSV Teaduste Akadeemia Toimetised. Füüsika-Matemaatika. Izvestiya Akademii Nauk Èstonskoĭ SSR. Seriya Fizika-Matematika, 35(3):338– 340, 344, 1986. CODEN ETFMB3. ISSN 0367-1429.

Mason:1970:OPA

[Mas70] J. C. Mason. Orthogonal polynomial approximation methods in numerical analysis. In Approximation Theory (Proceedings of the Symposium, Lancaster, 1969), pages 7–33. Academic Press, London-New York, 1970.

Matthews:1972:AOB

[Mat72] J. W. Matthews. An algorithm for obtaining best approximations on sets of positive polynomials. SIAM Journal on Numerical Analysis, 9(4):734–742, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/links.jstor.org/sici?sici=
0036-1429(197212)9:4<734:AAFOBA>
2.0.CO;2-T;   https://www.
jstor.org/stable/2156224.

Mason:2005:LPA

[MC05a] J. C. Mason and A. Crampton.

Laurent–Padé approximants to four kinds of Chebyshev polynomial expansions. II. Clenshaw– Lord type approximants. Numerical Algorithms, 38(1–3):19– 29, March 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mason:2005:LAAb

[MC05b] J. C. Mason and A. Crampton. Laurent–Padé approximants to four kinds of Chebyshev polynomial expansions. Part II: Clenshaw–Lord type approximants. Numerical Algorithms, 38 (1–3):19–29, March 2005. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Majidian:2022:SFC

[MFA22] Hassan Majidian, Medina Firouzi, and Amjad Alipanah. On the stability of Filon–Clenshaw–Curtis rules. Bulletin of the Iranian Mathematical Society, 48 (6):2943–2964, 2022. ISSN 1017-060X.

Malenge:1972:AMQ

[MP72] Jean-Pierre Malengé and Jean Pouget. Approximation en moyenne quadratique d’une fonction par une somme d’exponentielles. Cas du meilleur conditionnement. (French) [Mean square approximation of a function by a sum of exponentials. Case of the best packaging]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 275:A1347–A1349, 1972. CODEN CHASAP. ISSN 0151-0509.

Murty:1971:BAC

[Mur71] V. N. Murty. Best approximation with Chebyshev polynomials. SIAM Journal on Numerical Analysis, 8(4):717–721, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mason:1997:IMC

[MV97] J. C. Mason and E. Venturino. Integration methods of Clenshaw–Curtis type, based on four kinds of Chebyshev polynomials. In Günther Nürnberger, Jochen W. Schmidt, and Guido Walz, editors, Multivariate Approximation and Splines (Conference in Mannheim, September 7–10, 1996), volume 125 of Internat. Ser. Numer. Math., pages 153–165. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 3-7643-5654-5.

Neuman:1974:PUP

[Neu74]  E. Neuman.  Projections in

uniform polynomial approximations. Zastosowania Matematyki, 14:99–125, 1974. CODEN ZAM-TAK. ISSN 0044-1899.

Newbery:1974:EAP

[New74] A. C. R. Newbery. Error analysis for polynomial evaluation. Mathematics of Computation, 28 (127):789–793, July 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nikolajevic:2003:CFI

[NF03] Vladimir Nikolajevic and Gerhard Fettweis. Computation of forward and inverse MDCT using Clenshaw’s recurrence formula. IEEE Transactions on Signal Processing, 51(5):1439–1444, May 2003. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Notaris:2016:ENC

[Not16] Sotirios E. Notaris. The error norm of Clenshaw–Curtis and related quadrature formulae. BIT Numerical Mathematics, 56(2):705–728, June 2016. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

/link.springer.com/article/
10.1007/s10543-015-0569-6.

Novak:1996:HDI

[NR96] Erich Novak and Klaus Ritter. High-dimensional integration of smooth functions over cubes. Numerische Mathematik, 75(1):79– 97, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/00211/
bibs/6075001/60750079.htm;
 http://science.springer.de/
nmee/bibs/6075001/60750079.
htm.

Nobile:2015:CCC

[NTT15] Fabio Nobile, Lorenzo Tamellini, and Raul Tempone. Comparison of Clenshaw–Curtis and Leja quasi-optimal sparse grids for the approximation of random PDEs. In Spectral and High Order Methods for Partial Differential Equations—ICOSAHOM 2014, volume 106 of Lect. Notes Comput. Sci. Eng., pages 475–482. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-19800-9, 3-319-19799-1.

Oliver:1969:EET

[Oli69] J. Oliver. An error estimation technique for the solution of ordinary differential equations in Chebyshev series. The Computer Journal, 12(1):57–62, February 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Oliver:1971:PSC

[Oli71] J. Oliver. A practical strategy for the Clenshaw–Curtis quadrature method. Journal of the Institute of Mathematics and its Applications, 8(1):53–56, August 1971. CODEN JMTAA8. ISSN 0020-2932. URL https:

//academic.oup.com/imamat/
article-abstract/8/1/53/656064.

Oliver:1972:DAC

[Oli72] J. Oliver. A doubly-adaptive Clenshaw–Curtis quadrature method. The Computer Journal, 15(2): 141–147, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
15/2/141.full.pdf+html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/150141.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
02/tiff/141.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/
142.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_15/Issue_02/tiff/143.
tif;   http://www3.oup.co.

                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_02/tiff/144.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_15/Issue_
                                        02/tiff/145.tif;   http://
                                        www3.oup.co.uk/computer_journal/
                                        hdb/Volume_15/Issue_02/tiff/
                                        146.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_15/Issue_02/tiff/147.
                                        tif.

Oliver:1975:SME

[Oli75] J. Oliver. Stable methods for evaluating the points cos(iπ/n). Journal of the Institute of Mathematics and its Applications, 16 (2):247–257, October 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Oliver:1977:EAM

[Oli77] J. Oliver. An error analysis of the modified Clenshaw method for evaluating Chebyshev and Fourier series. Journal of the Institute of Mathematics and its Applications, 20(3):379–391, November 1977. CODEN JM-TAA8. ISSN 0020-2932.

Oliver:1980:ANT

[Oli80] J. Oliver. On the application of Newbery’s transformation to the Reinsch polynomial evaluation scheme. Journal of Computational and Applied Mathematics, 6(1):43–52, March 1980. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Olver:1967:BSS

[Olv67] F. W. J. Olver. Bounds for the solutions of second-order linear difference equations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 71B(4):161–166, October/December 1967. CO-DEN JNBBAU. ISSN 0022-4340 (print), 2376-5283 (electronic).

Olver:2004:CCM

[Olv04] Frank Olver. Charles Clenshaw: Mathematician with a flair for numerical analysis. The Guardian, ??(??):??, October 28, 2004. ISSN 0261-3077 (print), 1756-3224 (elec-

tronic). URL https://www.
theguardian.com/news/2004/
oct/29/guardianobituaries.
obituaries.

Ostrowski:1978:IRN

[OR78] Alexander M. Ostrowski and Raymond M. Redheffer. Inequalities related to the normal law. In General inequalities (Proceedings of the First International Conference, Mathematics Research Institute, Oberwolfach, 1976), I, volume 41 of Internat. Ser. Numer. Math., pages 125–129. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1978. ISBN 3-7643-0972-5.

Orlov:1977:ADE

[Orl77a] K. Orlov. Application of differential equations for transforming the implicit functions into the explicit ones. Matematički Vesnik, 1/14/29(3):295–302, 1977. ISSN 0025-5165 (print), 2406-0682 (electronic).

Orlov:1977:EQS

[Orl77b] K. Orlov. The elimination of quantities and the solving of systems of transcendent equations by means of differential equations. Matematički Vesnik, 1/14/29(3):303–308, 1977. ISSN 0025-5165 (print), 2406-0682 (electronic).

OHara:1968:EEC

[OS68] H. O’Hara and Francis J. Smith. Error estimation in the Clenshaw–Curtis quadrature formula. The Computer Journal, 11(2):213–219, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/
content/11/2/213.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_02/110213.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/213.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/214.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/
215.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/216.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/217.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
02/tiff/218.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_11/Issue_02/tiff/
219.tif.

Patent:1976:EQE

[Pat76] Paul D. Patent. The effect of quadrature errors in the computation of L^{2} piecewise polynomial approximations. SIAM Journal on Numerical Analysis, 13(3):344–361, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Piessens:1983:MCC

[PB83] Robert Piessens and Maria Branders. Modified Clenshaw–Curtis method for the computation of Bessel function integrals. BIT (Nordisk tidskrift for informationsbehandling), 23(3):370–381, September 1983. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=23&issue=3&spage=370.

Pati:2012:MQRb

[PDP12a] Anasuya Pati, Rajani B. Dash, and Pritikanta Patra. A mixed quadrature rule by blending Clenshaw–Curtis and Gauss– Legendre quadrature rules for approximate evaluation of real definite integrals. International Journal of Pure Mathematical Sciences, 2:10–15, June 2012. ISSN 2297-6205.

Pati:2012:MQRa

[PDP12b] Anasuya Pati, Rajani B. Dash, and Pritikanta Patra. A mixed quadrature rule by blending Clenshaw–Curtis and Lobatto quadrature rules for approximating real definite integrals. International Journal of Mathematical Sciences and Engineering Applications. IJMSEA, 6(2):423–430, 2012. ISSN 0973-9424.

Peherstorfer:1978:ESM

[Peh78] Franz Peherstorfer. Erweiterung des Satzes von Markoff. (German) [Extension of Markoff’s Theorem]. In Linear spaces and approximation (Proceedings of the Conference, Mathematics Research Institute, Oberwolfach, 1977), volume 40 of Internat. Ser. Numer. Math., pages 423–431. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1978. ISBN 3-7643-0979-2.

Petras:1994:HOE

[Pet94] Knut Petras. High order error constants of Clenshaw–Curtis product rules. Analysis, 14(1):1– 18, March 1994. ISSN 0174-4747 (print), 2196-6753 (electronic).

Piessens:1987:MCC

[Pie87] R. Piessens. Modified Clenshaw– Curtis integration and applications to numerical computation of integral transforms. In Numerical Integration (Halifax, N.S., 1986), volume 203 of NATO Adv. Sci. Inst. Ser. C: Math. Phys. Sci., pages 35–51. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2514-4.

Pryce:1981:REA

[Pry81] J. D. Pryce. Round-off error analysis with fewer tears. Bulletin of the Institute of Mathematics and its Applications, 17(2–3):40– 47, 1981. CODEN IMTABW. ISSN 0905-5628.

Press:1992:NRFa

[PTVF92a] William H. Press, Saul A.

Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical recipes in FORTRAN: the art of scientific computing. Cambridge University Press, Cambridge, UK, second edition, October 1, 1992. ISBN 0-521-43064-X (book), 0-521-43721-0 (example book) 0-521-43717-2 (diskette), 0-521-43719-9 (diskette), 0-521-43716-4 (diskette). xxvi + 963 pp. LCCN QA76.73.C15 N865 1992. US$39.95; US$54.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
052143064X;   http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0521437164;
http://www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&isbn=
0521437172;   http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0521437199. In-

cludes MacIntosh disk.

Press:1992:RRC

[PTVF92b] William H. Press, Saul A.

Teukolsky, William T. Vetterling, and Brian P. Flannery. Recurrence relations and Clenshaw’s recurrence formula. In NR-1992 [PTVF92a], chapter 5.5, pages 172–178. ISBN 0-521-43064-X (book), 0-521-43721-0 (example book) 0-521-43717-2 (diskette), 0-521-43719-9 (diskette), 0-521-43716-4 (diskette). LCCN QA76.73.C15 N865 1992. US$39.95; US$54.95.

URL http://www.cbooks.com/
sqlnut/SP/search/gtsumt?source=
&isbn=052143064X;   http:/
/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0521437164;   http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0521437172;
http://www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&isbn=

0521437199. Includes MacIntosh disk.

Randazzo:1977:DFE

[Ran77] D. J. Randazzo. Data fits with exponential functions. In Numerical analysis (Proceedings of the Colloquium, Lausanne, 1976), volume 37 of Internat. Ser. Numer. Math., pages 77–

94. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1977. ISBN 3-7643-0939-3.

Reimer:1977:CSS

[Rei77] M. Reimer. Clenshaw sums in several variables. In Constructive theory of functions of several variables (Proceedings of a Conference, Mathematics Research Institute, Oberwolfach, 1976), volume 571 of Lecture Notes in Math., pages 176–185. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977.

Rice:1964:WA

[Ric64] John R. Rice. On the L_{∞} Walsh arrays for Γ(x) and Erfc(x). Mathematics of Computation, 18 (88):617–626, October 1964. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riess:1971:EECa

[Rie71] R. D. Riess. Error estimates for Clenshaw–Curtis quadrature: preliminary report. Notices of the American Mathematical Society, 18(5):790– 791, August 1971. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
journals/notices/197108/197108FullIssue.
pdf.

Riess:1971:EECb

[RJ71] R. D. Riess and L. W. Johnson.

Error estimates for Clenshaw– Curtis quadrature. Numerische Mathematik, 18(4):345–353, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Riess:1972:EEC

[RJ72] R. D. Riess and L. W. Johnson.

Error estimates for Clenshaw– Curtis quadrature. Numerische Mathematik, 18(4):345–353, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roulier:1970:PBC

[Rou70] John A. Roulier. Permissable bounds on the coefficients of approximating polynomials. Journal of Approximation Theory, 3: 117–122, 1970. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Rowland:1969:LDP

[Row69] John H. Rowland. On the location of the deviation points in Chebyshev approximation by polynomials. SIAM Journal on Numerical Analysis, 6(1):118– 126, March 1969. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ray:1963:CPO

[RP63] W. D. Ray and A. E. N. T. Pitman. Chebyshev polynomial and other new approximations to Mills’ ratio. Annals of Mathematical Statistics, 34(3): 892–902, September 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.
org/euclid.aoms/1177704012.

Rosman:1971:CAR

[RR71] Bernard H. Rosman and Peter D. Rosenbaum. Chebyshev approximation of regulated functions by unisolvent families. Journal of Approximation Theory, 4: 113–119, 1971. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Roulier:1972:APR

[RT72] John A. Roulier and Gerald D. Taylor. Approximation by polynomials with restricted ranges of their derivatives. Journal of Approximation Theory, 5:216– 227, 1972. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Raj:2007:FCI

[RV07] P. Ananth Raj and A. Venkataramana. Fast computation of inverse Krawtchouk moment transform using Clenshaw’s recurrence formula. In 2007 IEEE International Conference on Image Processing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007.

Salzer:1973:RSC

[Sal73] Herbert E. Salzer. A recurrence scheme for converting from one orthogonal expansion into another. Communications of the ACM, 16(11):705–707, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saravi:2008:NSL

[SAM08] M. Saravi, F. Ashrafi, and S. R.

Mirrajei. Numerical solution of linear ordinary differential equations in quantum chemistry by Clenshaw method. International Journal of Computational and Mathematical Sciences, 2(4):186– 189, 2008. ISSN 2070-3910.

Schoenberg:1978:CMT

[Sch78] I. J. Schoenberg. On Chebyshev and Markov-type problems for polynomials in a circular ring. In Fourier analysis and approximation theory (Proceedings of the Colloquium, Budapest, 1976), Vol. II, volume 19 of Colloq. Math. Soc. János Bolyai, pages 697–712. North-Holland, Amsterdam, The Netherlands, 1978. ISBN 0-444-85197-6.

Schaback:1979:SEA

[Sch79]  Robert Schaback.  Subopti-

mal exponential approximations. SIAM Journal on Numerical Analysis, 16(6):1007–1018, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:2011:NEE

[SH11] H. V. Smith and D. B. Hunter. The numerical evaluation of the error term in a quadrature formula of Clenshaw–Curtis type for the Gegenbauer weight function. BIT Numerical Mathematics, 51(4):1031–1038, December 2011. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=51&issue=4&spage=1031.

Sugiura:2021:TCC

[SH21] Hiroshi Sugiura and Takemitsu Hasegawa. A truncated Clenshaw– Curtis formula approximates integrals over a semi-infinite interval. Numerical Algorithms, 86 (2):659–674, February 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://link.springer.
com/article/10.1007/s11075-
020-00905-w.

Sugiura:2022:ECC

[SH22] Hiroshi Sugiura and Takemitsu Hasegawa. Extensions of Clenshaw–Curtis-type rules to integrals over a semi-infinite interval. Numerical Algorithms, 90(1):3–30, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL https://link.springer.
com/article/10.1007/s11075-
021-01177-8.

Skandari:2020:DMB

[SHN20] Mohammad Hadi Noori Skandari, Marzieh Habibli, and Alireza Nazemi. A direct method based on the Clenshaw–Curtis formula for fractional optimal control problems. Mathematical Control and Related Fields, 10 (1):171–187, March 2020. ISSN 2156-8472 (print), 2156-8499 (electronic).

Smarzewski:1985:SOD

[Sma85] Ryszard Smarzewski. X-splines of odd degree. Journal of Approximation Theory, 44(4):295– 314, 1985. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Smith:1965:QMB

[Smi65] Francis J. Smith. Quadrature methods based on the Euler– Maclaurin formula and on the Clenshaw–Curtis method of integration. Numerische Mathematik, 7(5):406–411, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smith:1982:GEB

[Smi82]  H. V. Smith.  Global er-

ror bounds for the Clenshaw– Curtis quadrature formula. BIT (Nordisk tidskrift for informationsbehandling), 22(3):395–398, September 1982. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0006-3835&
                                        volume=22&issue=3&spage=395.

Smith:2015:CPQ

[Smi15]  H. V. Smith.  Convergence

properties of a quadrature formula of Clenshaw–Curtis type for the Gegenbauer weight function. BIT Numerical Mathematics, 55 (3):823–842, September 2015. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:/

                                        /link.springer.com/article/
                                        10.1007/s10543-014-0520-2.

Smoktunowicz:2002:BSC

[Smo02] Alicja Smoktunowicz. Backward stability of Clenshaw’s algorithm. BIT Numerical Mathematics, 42(3):600–610, September 2002. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0006-3835&
                                        volume=42&issue=3&spage=600.

Sommariva:2013:FCF

[Som13] Alvise Sommariva. Fast construction of Fejér and Clenshaw– Curtis rules for general weight functions. Computers and Mathematics and Applications, 65(4): 682–693, February 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S089812211200689X.

Sloan:1978:PCR

[SS78a] Ian H. Sloan and W. E. Smith.

Product-integration with the Clenshaw–Curtis and related points-convergence properties. Numerische Mathematik, 30(4): 415–428, December 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1978:PIC

[SS78b] Ian H. Sloan and W. E. Smith.

Product-integration with the Clenshaw–Curtis and related points. Convergence properties. Numerische Mathematik, 30(4): 415–428, December 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1980:PIC

[SS80] Ian H. Sloan and William E.

Smith. Product integration with the Clenshaw–Curtis points: implementation and error estimates. Numerische Mathematik, 34(4):387–401, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schumaker:1969:APH

[ST69] L. L. Schumaker and G. D. Taylor. On approximations by polynomials having restricted ranges. II. SIAM Journal on Numerical Analysis, 6(1):31–36, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stiefel:1965:PMP

[Sti65]  E. Stiefel.  Phase methods

for polynomial approximation. In Approximation of Functions (Proceedings of the Symposium at General Motors Research Laboratory, 1964), pages 68–82. Elsevier, Amsterdam, The Netherlands, 1965.

Sutton:2019:AAC

[Sut19a] Brian Sutton. Accuracy analysis for Clenshaw–Curtis quadrature. In Numerical Analysis: Theory and Experiments [Sut19c], chapter 18, pages 203–210. ISBN 1-61197-569-7 (print), 1-61197-570-0 (e-book). LCCN QA297 .S877 2019.

Sutton:2019:ICG

[Sut19b] Brian Sutton. Integration on a Chebyshev grid: Clenshaw– Curtis quadrature. In Numerical Analysis: Theory and Experiments [Sut19c], chapter 17, pages 195–201. ISBN 1-61197-569-7 (print), 1-61197-570-0 (ebook). LCCN QA297 .S877 2019.

Sutton:2019:NAT

[Sut19c] Brian Sutton. Numerical Analysis: Theory and Experiments. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, January 2019. ISBN 1-61197-569-7 (print), 1-61197-570-0 (e-book). xv + 431 pp. LCCN QA297 .S877 2019.

Singhal:1973:ASR

[SV73] K. Singhal and J. Vlach. Accuracy and speed of real and complex interpolation. Computing, 11(2):147–158, 1973. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sommariva:2008:NCC

[SVZ08] Alvise Sommariva, Marco Vianello, and Renato Zanovello. Nontensorial Clenshaw–Curtis cubature. Numerical Algorithms, 49(1–4): 409–427, December 2008. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=1017-1398&
volume=49&issue=1&spage=409.

Talbot:1976:UAP

[Tal76] A. Talbot. The uniform approximation of polynomials by polynomials of lower degree. Journal of Approximation Theory, 17(3): 254–279, 1976. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Taylor:1968:APH

[Tay68] G. D. Taylor. On approximation by polynomials having restricted ranges. SIAM Journal on Numerical Analysis, 5(2):258– 268, June 1968. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Taylor:1969:AFH

[Tay69] G. D. Taylor. Approximation by functions having restricted ranges. III. Journal of Mathematical Analysis and Applications, 27:241–248, 1969. CO-DEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Taylor:1970:OSA

[Tay70] G. D. Taylor. Optimal starting approximations for Newton’s method. Journal of Approximation Theory, 3:156–163, 1970. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Tuan:1972:CSE

[TE72] P. D. Tuan and David Elliott. Coefficients in series expansions for certain classes of functions. Mathematics of Computation, 26(117):213–232, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Todd:1967:ICZ

[Tod67]  John Todd.  Inequalities of

Chebyshev, Zolotareff, Cauer, and W. B. Jordan. In Inequalities (Proceedings of a Symposium at Wright-Patterson Air Force Base, Ohio, 1965), pages 321– 328. Academic Press, New York, USA, 1967.

Trefethen:2008:GQB

[Tre08] Lloyd N. Trefethen. Is Gauss quadrature better than Clenshaw– Curtis? SIAM Review, 50 (1):67–87, January 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.aip.
org/link/?SIR/50/67/1.

Tuan:1973:ECM

[Tua73] P. D. Tuan. An extension of Clenshaw’s method for linear differential equations. BIT (Nordisk tidskrift for informationsbehandling), 13(3):372–374, September 1973. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=13&issue=3&spage=372.

Turner:1989:NAP

[Tur89] Peter R. Turner, editor. Numerical Analysis and Parallel Processing: Lectures given at The Lancaster Numerical Analysis Summer School 1987, volume 1397 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51645-X. vi + 264 pp. Lectures from the Third S.E.R.C. Numerical Analysis Summer School held at the University of Lancaster, Lancaster, 1987.

Taylor:1974:EQP

[TW74] G. D. Taylor and J. Williams. Existence questions for the problem of Chebyshev approximation by interpolating rationals. Mathematics of Computation, 28(128): 1097–1103, October 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Urabe:1969:NSB

[Ura69] Minoru Urabe. Numerical solution of boundary value problems in Chebyshev series — a method of computation and error estimation. In Conference on Numerical Solution of Differential Equations (Dundee, 1969), volume 109 of Lecture Notes in Math., pages 40–86. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969.

Venkataramana:2011:RCF

[VR11] A. Venkataramana and P. Ananth Raj. Recursive computation of forward Krawtchouk moment transform using Clenshaw’s recurrence formula. In 2011 Third National Conference on Computer Vision, Pattern Recognition, Image Processing and Graphics. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 2011.

vanWyk:2019:CCT

[vWGB19] Hans-Werner van Wyk, Max Gunzburger, and John Burkardt. Clenshaw–Curtis type rules for statistics. SIAM Journal on Scientific Computing, ??(??): 1–26, ???? 2019. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL https://people.sc.fsu.
edu/~jburkardt/publications/
vgb_2016.pdf.

W:1963:RDT

[W.63] J. W. W. Reviews and descriptions of tables and books: 1 [C, D, E, L, X]. — C. W. Clenshaw, Chebyshev Series for Mathematical Functions, Mathematical Tables, v. 5, National Physics Laboratory, Her Majesty’s Stationery Office, London, 1962, iv + 36 p., 27.5 cm. Mathematics of Computation, 17(81):88, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Waldvogel:2006:FCF

[Wal06] Jörg Waldvogel. Fast construction of the Fejér and Clenshaw– Curtis quadrature rules. BIT Numerical Mathematics, 46(1): 195–202, March 2006. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0006-3835&
volume=46&issue=1&spage=195.

Wang:1992:ISS

[Wan92] Paul S. Wang, editor. International System Symposium on Symbolic and Algebraic Computation 92, July 27–29, Berkeley, California. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-489-9. LCCN QA 76.95 I59 1992.

Wang:2018:CRC

[Wan18] Haiyong Wang. On the convergence rate of Clenshaw–Curtis quadrature for integrals with algebraic endpoint singularities. Journal of Computational and Applied Mathematics, 333(??): 87–98, May 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042717305368.

Wills:1982:RCA

[WBR82] C. A. Wills, J. M. Blair, and P. L. Ragde. Rational Chebyshev approximations for the Bessel functions J_{0}(x), J_{1}(x), Y_{0}(x), Y_{1}(x). Mathematics of Computation, 39(160):617–623, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weideman:2007:KPF

[WT07] J. A. C. Weideman and L. N. Trefethen. The kink phenomenon in Fejér and Clenshaw–Curtis quadrature. Numerische Mathematik, 107(4):707–727, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Xiang:2012:CRG

[XB12] Shuhuang Xiang and Folkmar Bornemann. On the convergence rates of Gauss and Clenshaw– Curtis quadrature for functions of limited regularity. SIAM Journal on Numerical Analysis, 50(5):2581–2587, January 2012. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xiang:2013:CRG

[XB13] Shuhuang Xiang and Folkmar Bornemann. On the convergence rates of Gauss and Clenshaw– Curtis quadrature for functions of limited regularity. SIAM Journal on Numerical Analysis, 50(5):2581–2587, January 2013. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [XCWB13].

Xiang:2011:CCF

[XCWB11] Shuhuang Xiang, Yeol Je Cho, Haiyong Wang, and Hermann Brunner. Clenshaw–Curtis– Filon-type methods for highly oscillatory Bessel transforms and applications. IMA Journal of Numerical Analysis, 31(4):1281– 1314, October 2011. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Xiang:2013:ECC

[XCWB13] Shuhuang Xiang, Yeol Je Cho, Haiyong Wang, and Hermann Brunner. Erratum to “Clenshaw– Curtis–Filon-type methods for highly oscillatory Bessel transforms and applications” (IMA Journal of Numerical Analysis (2011)31: 1281–1314) [MR2846756]. IMA Journal of Numerical Analysis, 33(4): 1480–1483, May 2013. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). See [XB13].

Xiang:2015:EBF

[XHC15] Shuhuang Xiang, Guo He, and Yeol Je Cho. On error bounds of Filon–Clenshaw–Curtis quadrature for highly oscillatory integrals. Advances in Computational Mathematics, 41(3):573– 597, September 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Xiang:2014:FSI

[XHW14] Shuhuang Xiang, Guo He, and Haiyong Wang. On fast and stable implementation of Clenshaw– Curtis and Fejér-type quadrature rules. Abstract and Applied Analysis, page Art. ID 436164, 2014. ISSN 1085-3375 (print), 1687-0409 (electronic).

Xuan:2011:CCR

[XrL11] Yan Xuan and Fu rong Lin. Clenshaw–Curtis-rational quadrature rule for Wiener–Hopf equations of the second kind. In Recent Advances in Scientific Computing and Matrix Analysis [proceedings from the International Workshop on Scientific Computing and Matrix Analysis held at the University of Macau, China, from the 28th to 30th of December 2009], pages 93–110. International Press, Somerville, MA, USA, 2011. ISBN 1-57146-202-3. LCCN N32-532 I61/2009.